perimeter of a rectangle is 66 units when length and breadth are increased by 7 units each resulting area is 540 units.find the dimensions of original rectangle
Answer by ankor@dixie-net.com(22740)

perimeter of a rectangle is 66 units
2L + 2W = 66
simplify, divide by 2
L + W = 33
L = (33-W)
:
when length and breadth are increased by 7 units each resulting area is 540 units.
(L+7) * (W+7) = 540
replace L with (33-W)
((33-W)+7) * (W+7) = 540
(-W+40) * (W+7) = 540
FOIL
-W^2 - 7W + 40W + 280 - 540 = 0
A quadratic equation
-W^2 + 33W - 260 = = 0
multiply equation by -1
W^2 - 33w + 260 = 0
Factors to
(W-13)(W-20) = 0
Two solutions
W = 13 is the width, then L = 20
and
W = 26 is the width, then L = 13


find the dimensions of original rectangle: 20 by 13